# FreeStyle: Free Lunch for Text-guided Style Transfer using Diffusion Models

### v1

@@ 4.1 Implementation Details
Since our method is training-free, our method requires no training. Our experiments are conducted on an NVIDIA A100 GPU, with an average sampling time of about 31 seconds for a single image of $1024\times 1024$. All experiments in this section were completed with fixed hyperparameters, set as $n=320$, $s=1$, $b=2.5$, and $\sigma=958$. We use the DDIM sampler to execute a total of 30 sampling steps for each image generation. Our model is based on the SDXL Podell et al. (2023), utilizing its publicly available pre-trained model as the model parameters for inference.
@@ 4.2 Experimental Result
Qualitative Results. To verify the robustness and generalization ability of FreeStyle, we conducted numerous style transfer experiments with various styles across different content. Figure 3 presents the style transfer effects of FreeStyle in the domains of buildings, landscapes, animals, etc. The experiments include style transfer in “Chinese Ink”, “Embroidery Art”, “Oil Painting”, “Watercolor Painting”, “Studio Ghibli”, “Cyberpunk”, “Pixel Punk”, “Wasteland” and “Sketching”. We showcase the results of applying style transfer to human portraits using FreeStyle, as in Figure 4. In this figure, we also conducted style transfer experiments with multiple styles, including “Ufotable”, “Studio Ghibli”, “JOJO” and “Illumination Entertainment”. Observations indicate that FreeStyle is capable of providing accurate style information for the style transfer results while almost completely preserving the content information. For instance, the stylized results for “JOJO” maintain the structural information, while reasonably adjusting the image according to the character traits in the “JOJO” anime, like bold outlines, strong lines, and vibrant coloring. This achieves a more natural fusion and expression of both style and content. It is noteworthy that we performed style transfer on images using fine-grained styles from four animation categories in Figure 4. Despite this, FreeStyle is still able to achieve style transfer results with high recognizability and accurate styling.
[…]
@@
Quantitative Comparisons. To better evaluate our method, we employed multiple quantitative metrics for assessment, the results of which are presented in Table 1. For all comparison methods, we utilized their publicly available pretrained parameters for sampling. We performed style transfers on 202 content images including landscapes, architecture, people, and animals, across 10 styles (“Chinese Ink”, “Illumination Entertainment”, “Embroidery Art”, “Graffiti Art”, “Impressionism”, “Oil Painting”, “Watercolor Painting”, “Cyberpunk”, “Studio Ghibli”, “Sketching”), resulting in a total of 2020 stylized images for each method. For the CLIP Score Radford et al. (2021), we calculate the cosine similarity between the CLIP image embeddings and the prompt text embeddings. Using the prompt as a style description, we believe that a higher CLIP Score indicates a more accurate expression of style. The CLIP Aesthetic Score evaluates the quality, aesthetics, and artistic nature of images using a publicly available pre-trained art scoring model. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing.
CLIP Aesthetic Score $\uparrow$ CLIP Score $\uparrow$ Preference $\uparrow$
@@
FreeStyle (ours) 6.3148 25.615 54.6%
Table 1: Quantitative comparisons with state-of-the-art methods are conducted, using CLIP Aesthetic Score, CLIP Score, and human preferences as our evaluation criteria.
User Study. Additionally, a user study is conducted to evaluate the alignment of stylized images from six methods with human preferences. Forty participants from diverse fields and age groups are invited to evaluate 80 sets of different images, each transferred in various styles using 6 different methods. Participants are asked to select the best stylized image based on considerations of style expression, content preservation, and aesthetics. The “Preference” column in Table 1 reflects the percentage of instances where FreeStyle is ranked first by evaluators, showing its clear superiority over other methods.
@@ 4.3 Ablation Study
Effect of hyperparameters $b$ and $s$. We present the results of ablation experiments conducted on hyperparameters $b$
@@
and $s$, as in Figure 6. In FreeStyle, the intensities of content and style information are adjusted by the two hyperparameters $b$ and $s$, respectively. We observed that when the value of $b$ is smaller, the content in the image is more severely disturbed by the style. On the other hand, a larger value of $s$ leads to smoother image textures, while a very small $s$ value creates textures resembling noise. Overall, FreeStyle has relatively low sensitivity to hyperparameters, demonstrating strong robustness. Specifically, we find that its performance is optimal when the hyperparameter $b$ is set to 2.5 and $s$ is set to 1.
Effect of hyperparameter $\sigma$. Figure 7 illustrates the impact of the hyperparameter $\sigma$ on the style transfer effect. The observations indicate that better style transfer are achieved when $\sigma$ exceeds 850, whereas the effect gradually deteriorates as $\sigma$ becomes too small. We believe that a too small $\sigma$ value results in $f_{s}$ containing excessive content information, which significantly disrupts the style information.
@@
Figure 7: The ablation study of hyper-parameter $\sigma$.
@@ 5 Conclusion
In this work, we introduce FreeStyle, an innovative text-guided style transfer method leveraging pre-trained large text-guided diffusion models. Unlike prior methods, FreeStyle achieves style transfer without additional optimization or the need for reference style images. The framework, featuring a dual-stream encoder and a single-stream decoder, adapts seamlessly to specific style transfer through the adjustment of scaling factors. Despite its extreme simplicity, our method demonstrates superior performance in visual quality, artistic consistency and robust content information preservation across various styles and content images. These results contribute significantly to advancements in the field of training-free style transfer.

### v2

@@ 4.1 Implementation Details
Since our method is training-free, our method requires no training. Our experiments are conducted on an NVIDIA A100 GPU, with an average sampling time of about 31 seconds for a single image of $1024\times 1024$. As a training-free model, FreeStyle inevitably requires appropriate adjustment of hyperparameters to balance the intensity of style and content. In our qualitative experiments, we set the hyperparameters with $n=160$, $\sigma=958$, $b\in\left(0.5,3\right)$, and $s\in\left(0.5,2.5\right)$. We use the DDIM sampler to execute a total of 30 sampling steps for each image generation. Our model is based on the SDXL [4], utilizing its publicly available pre-trained model as the model parameters for inference.
Figure 4: The results of style transfer on portraits (left) and objects(right) using FreeStyle. Under the conditions of fine-grained anime style (left) and physical style(right) text prompts, the stylized results achieved with FreeStyle still exhibit clear fine-grained style differences and achieve a natural fusion of style and content.
Figure 5: Qualitative comparison with several state-of-the-art image style transfer methods, e.g., CLIPstyler [17], UDT2I [20], CAST [18], and StyTr2 [19]. For the “SDXL*” approach, we initially add $\delta$ ($\delta$ is set to 850) steps of noise to the content image, then guide the denoising for $T$ steps using style prompts to generate a stylized image.
@@ 4.2 Experimental Result
Qualitative Results. To verify the robustness and generalization ability of FreeStyle, we conduct numerous style transfer experiments with various styles across different content. Figure 3 presents the style transfer effects of FreeStyle in the domains of buildings, landscapes, animals, etc. The experiments include style transfer in “Chinese Ink”, “Embroidery Art”, “Oil Painting”, “Watercolor Painting”, “Studio Ghibli”, “Cyberpunk”, “Pixel Punk”, “Wasteland” and “Sketching”. We showcase the results of applying style transfer to human portraits using FreeStyle, as in Figure 4 (left). In this figure, we also conduct style transfer experiments with multiple styles, including “Ufotable”, “Studio Ghibli”, “JOJO” and “Illumination Entertainment”. Observations indicate that FreeStyle is capable of providing accurate style information for the style transfer results while almost completely preserving the content information. For instance, the stylized results for “JOJO” maintain the structural information, while reasonably adjusting the image according to the character traits in the “JOJO” anime, like bold outlines, strong lines, and vibrant coloring. This achieves a more natural fusion and expression of both style and content. It is noteworthy that we perform style transfer on images using fine-grained styles from four animation categories in Figure 4 (left). Despite this, FreeStyle is still able to achieve style transfer results with high recognizability and accurate styling. Additionally, we apply multiple physical style transfers to various everyday items. As illustrated in Figure 4 (right), FreeStyle demonstrates excellent style transfer effects across these styles.
Qualitative Comparisons. As shown in Figure 5, we conduct extensive comparative experiments with state-of-the-art methods, covering various styles and diverse content images. The results of “SDXL*” (e.g., rows 3,5) are highly sensitive to its hyperparameter $\delta$, resulting in significant variance in its stylized images. In contrast, FreeStyle achieves effective style transfer more consistently, demonstrating stronger robustness for images of varying content and across diverse style transfer tasks. The results show the apparent advantages of our method over others, as it can reasonably modify shapes (e.g., rows 1,2,6), brushstrokes (e.g., rows 1-5), lines (e.g., rows 3,4), and colors (e.g., rows 1-6) to achieve superior artistic effects. In comparisons between our method and several others, it is noticeable that our approach more accurately achieves style expression (e.g., rows 2,3,6), especially in styles that are more challenging to transfer. Results from CAST [18] and StyTr2 [19] are often marked by noticeable halo effects (e.g., rows 3,5,6) and are blurred (e.g., rows 2,6). In contrast, FreeStyle can produce clear stylized images without any noticeable halo effects. In summary, Figure 5 indicates that our method exhibits greater robustness, more accurate style expression, and more artistic style transfer effects.
@@
Quantitative Comparisons. To better evaluate our method, we employed multiple quantitative metrics for assessment, the results of which are presented in Table 1. For all comparison methods, we utilized their publicly available pretrained parameters for sampling. Following the widely used quantitative experimental setup [7, 17], we performed style transfers on 202 content images including landscapes, architecture, people, and animals, across 10 styles (“Chinese Ink”, “Illumination Entertainment”, “Embroidery Art”, “Graffiti Art”, “Impressionism”, “Oil Painting”, “Watercolor Painting”, “Cyberpunk”, “Studio Ghibli”, “Sketching”), resulting in a total of 2020 stylized images for each method. For the CLIP Score [37], we calculate the cosine similarity between the CLIP image embeddings and the prompt text embeddings. Using the prompt as a style description, we believe that a higher CLIP Score indicates a more accurate expression of style. The CLIP Aesthetic Score evaluates the quality, aesthetics, and artistic nature of images using a publicly available pre-trained art scoring model. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing. Training Cost refers to the product of the number of parameters that need to be optimized during the training phase and the recommended number of iterations in the corresponding method.
@@
FreeStyle (ours) 6.3148 25.615 0M 54.6%
Table 1: Quantitative comparisons with state-of-the-art methods are conducted, using CLIP Aesthetic Score, CLIP Score, Training Cost and human preferences as our evaluation criteria.
User Study. In addition to CLIP Aesthetic Score and CLIP score, we conduct a user study to evaluate the alignment of stylized images from six methods with human preferences. Forty participants from diverse fields and age groups are invited to evaluate 80 sets of different images, each transferred in various styles using 6 different methods. Participants are asked to select the best stylized image based on considerations of style expression, content preservation, and aesthetics. The “Preference” column in Table 1 reflects the percentage of instances where FreeStyle is ranked first by evaluators, showing its clear superiority over other methods.
@@ 4.3 Ablation Study
Effect of hyperparameters $b$ and $s$. We present the results of ablation experiments conducted on hyperparameters $b$ and $s$, as in Figure 6. In FreeStyle, the intensities of content and style information are adjusted by the two hyperparameters $b$ and $s$, respectively. We observed that when the value of $b$ is smaller, the content in the image is more severely disturbed by the style. On the other hand, a larger value of $s$ leads to smoother image textures, while a very small $s$ value creates textures resembling noise. Overall, FreeStyle shows low sensitivity to hyperparameters, demonstrating robustness.
@@
Effect of hyperparameters $s$. We conduct ablation experiments on the transfer of “origami art” style using different settings of $s$, in Figure 7. It is evident that adjusting the hyperparameter $s$ significantly affects the intensity of the style in the images. As $s$ increases, the style intensity enhances while the content information is relatively diminished. Conversely, reducing $s$ weakens the style intensity and can even lead to inaccuracies in style expression, as seen in the second row of Figure 7 where $s=0.2$.
Figure 7: Ablation experiment on the impact of the hyperparameter $s$ on style intensity.
Effect of hyperparameter $\sigma$. Figure 8 illustrates the impact of the hyperparameter $\sigma$ on the style transfer effect. The observations indicate that better style transfer are achieved when $\sigma$ exceeds 850, whereas the effect gradually deteriorates as $\sigma$ becomes too small. We believe a too small $\sigma$ value makes $f_{s}$ contain excessive content information, significantly disrupting the style information.
Figure 8: The ablation study of hyper-parameter $\sigma$.
@@ 5 Conclusion
In this study, we present FreeStyle, an innovative text-guided style transfer method that utilizes pre-trained large text-guided diffusion models. Diverging from previous approaches, FreeStyle accomplishes style transfer without the need for additional optimization or reference style images. The framework, comprising a dual-stream encoder and a single-stream decoder, seamlessly adapts to specific style transfer tasks by adjusting scaling factors. Despite its simplicity, our method showcases superior performance in terms of visual quality, artistic consistency, and robust preservation of content information across diverse styles and content images. These findings significantly advance the field of training-free style transfer.
A limitation of our algorithms is that for some challenging content images or style types, frequent manual adjustment of hyperparameters is required. However, we believe that with the diffusion model becoming more powerful and improvements in model structure, this issue will gradually diminish.